OpenStudy (anonymous):

classify the number as prime, composite oe neither 51 Also simplify 80/60

OpenStudy (mertsj):

Did you apply the divisibility tests to the number 51?

OpenStudy (anonymous):

no what is that

OpenStudy (mertsj):

Well, if it is even, it can be divided by 2. Is it an even number?

OpenStudy (anonymous):

it would be an odd number

OpenStudy (mertsj):

So it fails the divisible by 2 test. Add the digits. What do you get?

OpenStudy (mertsj):

Hello. Are you there?

OpenStudy (anonymous):

add which digits

OpenStudy (anonymous):

sorry I am new to algebra

OpenStudy (mertsj):

The digits of the number 51

OpenStudy (mertsj):

5+1=???

OpenStudy (anonymous):

you get 6

OpenStudy (anonymous):

51 must be prime i suppose

OpenStudy (mertsj):

So if the sum of the digits is divisible by 3, then the number is divisible by 3. If it is divisible by an integer, then it is NOT prime and is composite.

OpenStudy (anonymous):

so I would classify 51 as neither prime or composite

OpenStudy (anonymous):

51= 17 x 3 so it is composite

OpenStudy (anonymous):

80/60 is 8/6 which is 4/3 or 1 1/3

OpenStudy (anonymous):

51 is divisible by 3, therefore it is a composite number. Prime numbers are those that cannot be divided by anything except for 1 and the number itself. Composite numbers have other factors. Neither usually indicates that the number is complex or imaginary (or 0). 80 / 60 = 8 / 6 = 4 / 3 = 1 and 1/3

OpenStudy (anonymous):

Thanks to both of you for answers